OpenStudy (anonymous):

If the dimensions of a rectangle become one-fifth of the original dimensions, then area of the new rectangle will be _______of the area of the original rectangle.

OpenStudy (anonymous):

@saifoo.khan do you know how to find this? :)

OpenStudy (saifoo.khan):

That's cake.

OpenStudy (anonymous):

Can I use any numbers, as long as it's a fifth?

OpenStudy (saifoo.khan):

that's a general formula. But let's break it down.

OpenStudy (saifoo.khan):

Hold on.

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

I don't understand the formula. Can you try to explain it? I'm sorry

OpenStudy (saifoo.khan):

It's okay. i know it's confusing coz i modified it. hehe

OpenStudy (saifoo.khan):

Let me find out the easier way out. ;)

OpenStudy (saifoo.khan):

Got it. :D

OpenStudy (saifoo.khan):

\[A = (L)^2\]Simplified version. ;)

OpenStudy (anonymous):

Okay, what is A and L? Are there specific numbers?

OpenStudy (anonymous):

I mean, what do they stand for?

OpenStudy (saifoo.khan):

P.S. sorry i forgot to mention them, A = area L = length. Now, your A is turned down into 1/5. so make A = 1/5 \[\frac15 A = (L)^2\]

OpenStudy (saifoo.khan):

Now solve for L.

OpenStudy (anonymous):

Would the answer be 25?

OpenStudy (saifoo.khan):

Nope.

OpenStudy (saifoo.khan):

take square roots on both sides, \[\sqrt{\frac{1}{5}} = L\] \[L = \frac{1}{\sqrt5}\] \[L = \frac{\sqrt{5}}{5}\]

OpenStudy (anonymous):

.44?

OpenStudy (saifoo.khan):

Nope. that's the answer.

OpenStudy (anonymous):

That is the answer?

OpenStudy (saifoo.khan):

i just read the question again and found out, i messed up the working. :( I'm sorry.

OpenStudy (saifoo.khan):

Let me redo.

OpenStudy (saifoo.khan):